Consider two random variables ? ∈ {0, 1, 2} and ? ∈ {1, 2} with marginal...

  1. Consider two random variables ? ∈ {0, 1, 2} and ? ∈ {1, 2} with marginal distributions ??(?) = 1 for ? = 0,1,2 ?? (?) = 1 for ? = 1,2

    (a) Come up with two di erent joint PMF-s that give these marginals
    (?*) Come up with a joint PMF that gives these marginals and guarantees that

    ?(? ?)=1
    (?*) Come up with a joint PMF that makes ? (? ̸= ? ) as small as possible
Answer #1

Let xeço, 1, 2} & y = $1,23 with. fx (2) = 3 x = 0, 1, 2 , foly) = 2 3 4 = 1,2 Let X&Y be independent .. fxymiy ] = fxxl fyly

سیمه - Sxy 11,Y) Z=x-Y) م о сох - we cod at o wt م - رN flund F-X = 2 TO 11-12 lž P(x + 4) = P(x-40) -P(2+0) - P(22)+1(2=-1)